OpenStudy (anonymous):

f(x)=√5-5x x=-1,6 evaluate

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

setting -1.6 = -16/10 = -8/5 then put x=-8/5 to f(x) so, it can be